OpenStudy (anonymous):

Use the elimination method to solve the given system of equations. 5x -3y=16 -5x +2y=-15 a.(3.32,0.2) b.(3.8,1) c.(2.6,-1) d.(3.4,-1)

OpenStudy (anonymous):

equation 1) \[5x -3y =16 \] equation 2) \[-5x +2y=-15\] in this case we do not need to do anything to the equations before we add them. do equation 1 + equation 2 \[5x -3y +(-5x +2y) =15 +16\] now gather together the terms and you will we lose all the x's allowing us to solve for y

OpenStudy (anonymous):

that last line should be 16 - 15

OpenStudy (anonymous):

\[5x−3y+(−5x+2y)=16 -15\]

OpenStudy (kropot72):

Adding the two equations and solving for y leads to the answer (2.6,-1)

OpenStudy (anonymous):

Thanks :)

OpenStudy (kropot72):

You're welcome:)